# m2mKD: Module-to-Module Knowledge Distillation for Modular Transformers

### v1

@@ 4.1 Setups
Datasets. In this paper, we focus on the image classification task. The training for both the preparation and m2mKD phases is conducted using the ImageNet-1k dataset (Deng et al.,, 2009). For the NAC models, we perform end-to-end training using both the ImageNet dataset and its miniature version, Tiny-ImageNet. For OOD evaluation, we utilize the ImageNet-R(enditions) dataset (Hendrycks et al., 2021b, ) and its down-sampled subset, Tiny-ImageNet-R. Additionally, the CIFAR-100 dataset (Krizhevsky,, 2009) is employed for few-shot adaptation.
Preparation. Since our target NAC model consists of a total of 10 layers ($L=10$), we choose the DeiT-Huge model (Dosovitskiy et al.,, 2021) with 32 layers as the teacher model to ensure sufficient depth. The DeiT-Huge model is divided into 10 sub-modules, with the first and last sub-modules containing 4 layers each, and the remaining sub-modules comprising 3 layers each. The NAC models contain no more than 37M parameters, while the DeiT-Huge model contains 632M parameters. This yields approximately 3.7M and 63.2M parameters for each student and teacher module, respectively. At the beginning of the preparation stage, a 10-layer meta model is trained on ImageNet for 300 epochs. Subsequently, each teacher module is incubated by the meta model for 100 epochs, and all modules are assembled back together for additional fine-tuning of 100 epochs. The training configurations for this stage are identical to the Deep Incubation approach.
@@
E2E training. When training on the same dataset, the hyperparameters remain unchanged for both the reproduced baselines and the distilled NAC models. However, since the original paper on NACs does not provide a comprehensive list of hyperparameters and we were unable to reproduce the reported results using the given hyperparameters, we adjusted some of them in order to approach their reported results as closely as possible. Table 9 presents a selection of our hyperparameters [^2] For the full list of hyperparamters, please refer to the configuration files provided on GitHub., including all altered values.
@@ 4.2 Results
Main results. As aforementioned, our teacher model DeiT-Huge and NAC student modules in the first two phases are trained solely on ImageNet. The assembled teacher model achieves a validation accuracy of 81.8%. With 8 A100 80GB GPUs, the training of m2mKD for all ten student modules can be completed within 12 hours. For E2E training on Tiny-ImageNet, a portion of the input tokenizer and the entire output tokenizer in the student modules need to be discarded due to dimension mismatching. Table 1 compares the reproduced baselines and our distilled NAC models on Tiny-ImageNet and Tiny-ImageNet-R. The reproduced results slightly outperform the reported values in the original paper. Although the dataset is different from the one used in the m2mKD phase, our distilled models with various graph prior regularizers exhibit an average improvement of 5.3% in IID performance and 3.8% in OOD robustness over the baselines. This indicates that the addition of distilled student modules not only enhances the ability of the final NAC for similar tasks, but also improves its modularity. To ensure reproducibility, we repeat the E2E training phase of the distilled model with scale-free prior three times and measure their mean and standard deviation. The results are presented in Table 2.
The comparison for ImageNet is shown in Table 3. The original paper only provides the validation accuracy of the NAC trained on ImageNet with a scale-free graph prior, which is 77%. We reproduce the baselines for all four graph priors, with the maximum value being 76.1%. Our distilled NACs achieve maximum gains of 1.0% and 2.4% for IID and OOD performance, respectively.
Few-shot. To evaluate the few-shot adaptation performance, we further fine-tune the classifier layer of the distilled NAC with a scale-free prior, which is trained on ImageNet, using a small number of samples from the CIFAR-100 dataset. The hyperparameters are listed in Table 10, and the results are presented in Table 4. We conduct the experiments with 5 different seeds and report the averaged accuracies and corresponding standard deviations in Figure 5. Our reproduced baselines are at most 10% higher than the original results. It can be observed that the distilled model performs similarly to the baseline with no significant improvement. The standard deviations are relatively large, possibly due to the limited number of repetitions (i.e., the number of seeds). We also rerun the 2-shot experiments once for all five seeds to examine the variation under the same seeds. Table 5 shows that given a fixed seed, the largest difference is around 7%. Therefore, additional experiments may be necessary to further validate the few-shot performance.
@@
(Tiny-ImageNet Validation) (Tiny-ImageNet-R)
Acc@1 Acc@5 Acc@1 Acc@5
NACs with Scale-Free Prior 60.83 82.35 20.74 41.71
with Planted-Partition Prior 60.57 82.20 20.50 42.94
with Ring-of-Cliques Prior 60.70 82.57 20.89 41.80
with Erdos-Renyi Prior 61.53 83.14 21.08 42.25
+m2mKD with Scale-Free Prior 66.47 (+5.64) 85.08 (+2.73) 24.31 (+3.57) 44.38 (+2.67)
with Planted-Partition Prior 66.04 (+5.47) 85.68 (+3.48) 24.69 (+4.19) 45.36 (+2.42)
with Ring-of-Cliques Prior 66.21 (+5.51) 85.49 (+2.92) 24.89 (+4.00) 45.74 (+3.94)
with Erdos-Renyi Prior 65.99 (+4.46) 85.37 (+2.23) 24.54 (+3.46) 45.45 (+3.20)
IID OOD
Acc@1 Acc@5 Acc@1 Acc@5
1st 66.47 85.08 24.31 44.38
2nd 66.13 85.33 24.84 45.00
3rd 65.94 85.49 25.29 45.61
MEAN 66.18 85.30 24.81 45.00
STDEV 0.27 0.21 0.49 0.62
@@
(ImageNet Validation) (ImageNet-R)
Acc@1 Acc@5 Acc@1 Acc@5
NACs with Scale-Free Prior 75.61 93.93 37.30 54.01
with Planted-Partition Prior 75.71 94.09 37.63 54.38
with Ring-of-Cliques Prior 76.12 94.35 37.21 53.70
with Erdos-Renyi Prior 75.71 93.95 36.48 53.39
+m2mKD with Scale-Free Prior 76.63 (+1.02) 94.62 (+0.69) 39.18 (+1.88) 55.10 (1.09)
with Planted-Partition Prior 76.49 (+0.78) 94.01 (-0.08) 38.02 (+0.39) 53.58 (-0.8)
with Ring-of-Cliques Prior 76.89 (+0.77) 94.44 (+0.09) 39.29 (+2.08) 55.42 (+1.72)
with Erdos-Renyi Prior 76.56 (+0.85) 94.39 (+0.44) 38.84 (+2.36) 54.69 (+1.3)
Seed 1-shot 2-shot 4-shot 8-shot
E2E m2mKD E2E m2mKD E2E m2mKD E2E m2mKD
1 53.66 60.82 64.39 64.39 71.21 75.06 81.03 80.53
2 60.82 50.08 64.39 59.03 75.06 78.91 84.04 82.03
3 67.97 67.97 75.13 67.97 78.91 79.87 80.53 79.03
4 60.82 60.82 67.97 57.24 75.06 78.91 84.04 82.53
5 50.08 60.82 60.82 60.82 73.14 78.91 80.03 79.53
MEAN 58.67 60.10 66.54 61.89 74.68 78.33 81.93 80.73
STDEV 6.98 6.40 5.43 4.31 2.85 1.88 1.95 1.52
Figure 5: Few-shot performance of NACs on CIFAR-100..
Table 5: Differences between two runs of the 2-shot experiment on NACs.
seed E2E m2mKD
1st 2nd $|\Delta|$ 1st 2nd $|\Delta|$
1 64.39 69.76 5.37 64.39 66.18 1.79
2 64.39 64.39 0.00 59.03 60.82 1.79
3 75.13 75.13 0.00 67.97 75.13 7.16
4 67.97 66.18 1.79 57.24 62.60 5.36
5 60.82 62.60 1.78 60.82 64.39 3.57
MEAN 66.54 67.61 1.07 61.89 65.82 3.93
STDEV 5.43 4.97 0.46 4.31 5.57 1.26
@@ 5 Experiments on MoEs
In a MoE model, some or all of the feedforward networks (FFNs) in a standard monolithic model are replaced with MoE layers. As illustrated in Figure 3(b), a MoE layer is constructed by multiple experts and a gate, where each expert is essentially an FFN. The gate, typically implemented as a MLP, is responsible for selecting a specified number of experts to process the input. The outputs of the selected experts are then aggregated. The computation performed by a MoE layer can be expressed as (Shazeer et al.,, 2017):
$y=\sum_{i=1}^{n}G(x)_{i}E_{i}(x)$
@@ 5.1 Setups
Datasets. We exclusively used the ImageNet-1k dataset for MoE models throughout the pipeline (from preparation to E2E training). The few-shot adaptation ability is evaluated on both CIFAR-100 and CUB-2011 datasets (Wah et al.,, 2011).
@@
m2mKD. In this series of experiments, the target model $S$ is set as a Vision MoE Base (V-MoE-B) model (Riquelme et al.,, 2021). Instead of using the carefully designed gate proposed by (Riquelme et al.,, 2021), we employ the earliest introduced gate in (Shazeer et al.,, 2017) which purely performs a top-$k$ ($k=2$) operation and no additional constraint or balancing loss is applied:
$Gate(x)=Softmax(TopK(W\cdot x)).$
@@
Unlike the NAC models, the student modules of MoE do not require modification, and all of their learned parameters can be loaded into the target model during the end-to-end training phase. The same as the settings for NAC models, we set the balancing factor $\alpha=0.5$ and softmax temperature $\tau=1.0$. All of the remaining hyperparameters are identical to those used in Deep Incubation for incubating DeiT-B (Touvron et al.,, 2021) and the student modules are trained for 100 epochs.
@@ 5.2 Results
Main results. We compare m2mKD with three baselines: pure end-to-end training, conventional knowledge distillation, and Deep Incubation. For end-to-end training, we train a V-MoE-B model from scratch for 300 epochs using the same hyperparameters as in the DeiT-B training (Touvron et al.,, 2021). In contrast to m2mKD, which performs knowledge distillation at the module level, conventional KD refers to knowledge distillation between complete models. For the sake of fair comparison, we again use DeiT-L as the teacher model and V-MoE-B as the student model. The KL divergence between their output logits is incorporated into the loss with $\alpha=0.5$ and $\tau=1.0$ (see Equation 3). The training process lasts for 300 epochs. Lastly, we consider the original Deep Incubation approach, where we use their open source checkpoint of the meta model, originally trained for incubating DeiT-B, for the V-MoE-B experiments. The validation results on ImageNet-1k are summarized in Table 6. Conventional KD underperforms the other methods, with an accuracy 1.9% lower than pure end-to-end training, which is the second lowest. This verifies our assertion that existing knowledge distillation techniques are not necessarily compatible with modular models. On the other hand, the Deep Incubation approach achieves 2.97% higher accuracy than end-to-end training, demonstrating its effectiveness not only for monolithic models, but also for modular models. Our m2mKD approach further increases the accuracy by 0.50%, resulting in the best performance among these methods. Given the remarkable OOD results of NACs in previous experiments, we also investigate the OOD robustness of MoEs on the ImageNet-R dataset, which is rare to be discussed in other MoE-related literatures. Surprisingly, all the MoE models trained using the four different methods achieve near 0% accuracy on the ImageNet-R dataset. Based on these experiments, it appears that NACs are significantly stronger than MoEs in terms of OOD or zero-shot performance.
@@
Training time. In addition to accuracy, Table 6 presents the ratios of training time calculated relative to the time taken by the pure end-to-end training approach. m2mKD is slightly slower than Deep Incubation but roughly equivalent to the pure end-to-end method in terms of time, while conventional KD requires 1.12$\times$ time. Further experiments (Appendix C) demonstrate that the ratio for m2mKD can be reduced to 0.57 with marginal performance degradation. For a fair comparison, we exclude the duration of teacher model training for the KD and m2mKD methods. In other words, only the time for meta model training, distillation, and end-to-end training phases are considered in the time calculation for m2mKD. At this point, it is natural to question whether the preparation phase of m2mKD can be simplified. The current m2mKD pipeline involves incubating a teacher model in the preparation phase. We suppose that such a model would be a more knowledgeable teacher specifically for modular students. However, since numerous pretrained models are publicly available, it would be time-saving to utilize them as the teacher model for m2mKD instead of training a new one from scratch. If this approach is adopted, the total training time would be exactly as stated in the table. We leave the investigation of how teachers trained using different methods influence the performance of m2mKD in future work.
@@
Few-shot. We evaluate the few-shot adaptation of three models: a DeiT-B model trained using Deep Incubation, and two V-MoE-B models trained using Deep Incubation and m2mKD, respectively. The hyperparameters can be found in Table 10. During the few-shot training, all parameters in these models are frozen, except for the newly random initialized classifier layer, whose output dimension is set to the number of few-shot classes (8 in all experiments). The experiments for each shot are repeated with five different seeds. In Table 7, we report the highest accuracy achieved within 100 epochs for each experiment, along with the mean and standard deviation calculated over the seeds. The averaged results are also depicted in Figure 6. Surprisingly, the DeiT-B model outperforms both V-MoE-B models, which are expected to be capable due to their modularity. This might be attributed to the relatively small dataset size (ImageNet-1k), resulting in insufficient training samples for each expert to learn comprehensively. However, when focusing on the two MoE models, the one trained using m2mKD consistently outperforms the Incubation model by approximately 2% for all shots on CUB-2011, as well as the 1-shot experiment on CIFAR-100. Similar to the few-shot results of NACs (Section 4.2), the standard deviations are relatively large. Conducting experiments on a larger dataset is necessary to obtain a more precise evaluation of the few-shot performance.
Table 7: Few-shot performance of DeiT and V-MoE-B models on CIFAR-100 (top) and CUB-2011 (bottom). The models are tasked with classifying samples into 8 classes (i.e., 8-way). We compare three models: a DeiT-B model trained using Deep Incubation (DeiT-Inc), and two V-MoE-B models trained using Deep Incubation (MoE-Inc) and m2mKD (MoE-m2mKD), respectively.
CIFAR-100
Seed 1-shot 5-shot 10-shot
DeiT-Inc MoE-Inc MoE-m2mKD DeiT-Inc MoE-Inc MoE-m2mKD DeiT-Inc MoE-Inc MoE-m2mKD
1 70.83 70.83 70.83 90.83 90.00 89.17 90.83 91.67 90.42
2 70.83 62.50 66.37 90.00 87.50 86.67 92.92 92.50 90.42
3 87.50 95.83 91.67 95.83 94.17 94.17 97.08 96.25 96.67
4 95.83 79.17 87.50 95.83 94.17 94.17 97.50 94.58 96.25
5 79.17 70.83 70.83 88.33 87.50 87.50 90.00 89.58 90.42
MEAN 80.83 75.83 77.44 92.16 90.67 90.34 93.67 92.92 92.84
STDEV 10.87 12.64 11.33 3.47 3.36 3.61 3.48 2.59 3.31
CUB-2011
Seed 1-shot 5-shot 10-shot
DeiT-Inc MoE-Inc MoE-m2mKD DeiT-Inc MoE-Inc MoE-m2mKD DeiT-Inc MoE-Inc MoE-m2mKD
1 83.33 70.83 70.83 90.83 90.83 91.67 95.00 94.38 93.75
2 79.17 83.33 87.5 94.17 93.33 94.17 98.75 95.63 98.13
3 95.83 91.67 95.83 95.00 92.50 92.50 94.38 91.25 91.25
4 87.50 75.00 79.17 88.33 80.83 84.17 84.38 80.63 86.25
5 66.67 70.83 75.00 90.83 88.33 94.17 96.25 93.75 95.63
MEAN 82.50 78.33 81.67 91.83 89.16 91.34 93.75 91.13 93.00
STDEV 10.78 9.04 10.03 2.73 5.04 4.15 5.50 6.08 4.54
(a) CIFAR-100 8 way
(b) CUB-2011 8 way
Figure 6: Few-shot performance of DeiT and V-MoE-B models on CIFAR-100 (left) and CUB-100 (right).
@@
Ablations. In the proposed m2mKD pipeline, the stitch layers are used only during the distillation phase and subsequently discarded to ensure no modifications are required for the student model. Although the stitch layers typically account for just a small proportion of the total parameters in a large modular model (e.g., 1% for our V-MoE-B student), directly removing them may result in the loss of some knowledge acquired from the teacher, potentially impacting the final model’s performance negatively. To investigate this aspect, we conduct additional experiments by preserving the stitch layers in the student model. These experiments include decreasing the number of distillation epochs for m2mKD and scaling up. For further details, please refer to Appendix C.

### v2

@@ 4.1 Setups
Datasets. In this paper, we focus on the image classification task. The training for both the preparation and m2mKD phases is conducted using the ImageNet-1k dataset [3]. For the NAC models, we perform end-to-end training using both the ImageNet dataset and its miniature version, Tiny-ImageNet. For OOD evaluation, we utilize the ImageNet-R(enditions) dataset [10] and its down-sampled subset, Tiny-ImageNet-R. Additionally, the CIFAR-100 dataset [18] is employed for few-shot adaptation.
Preparation. Since our target NAC model consists of a total of 10 layers ($L=10$), we choose the DeiT-Huge model [4] with 32 layers as the teacher model to ensure sufficient depth. The DeiT-Huge model is divided into 10 sub-modules, with the first and last sub-modules containing 4 layers each, and the remaining sub-modules comprising 3 layers each. The NAC models contain no more than 37M parameters, while the DeiT-Huge model contains 632M parameters. This yields approximately 3.7M and 63.2M parameters for each student and teacher module, respectively. At the beginning of the preparation stage, a 10-layer meta model is trained on ImageNet for 300 epochs. Subsequently, each teacher module is incubated by the meta model for 100 epochs, and all modules are assembled back together for additional fine-tuning of 100 epochs. The training configurations for this stage are identical to the Deep Incubation approach.
@@
E2E training. When training on the same dataset, the hyperparameters remain unchanged for both the reproduced baselines and the distilled NAC models. Since the original paper on NACs does not provide a comprehensive list of hyperparameters and we were unable to reproduce the reported results using the given hyperparameters, we adjusted some of them in order to approach their reported results as closely as possible. Appendix B presents a selection of our hyperparameters, including all altered values.
@@ 4.2 Results
(Tiny-ImageNet) (Tiny-ImageNet-R)
Graph prior Acc@1 Acc@5 Acc@1 Acc@5
NACs Scale-Free 60.83 82.35 20.74 41.71
Planted-Partition 60.57 82.20 20.50 42.94
Ring-of-Cliques 60.70 82.57 20.89 41.80
Erdos-Renyi 61.53 83.14 21.08 42.25
[…]
Planted-Partition 66.04 $\uparrow$5.47 85.68 $\uparrow$3.48 24.69 $\uparrow$4.19 45.36 $\uparrow$2.42
Ring-of-Cliques 66.21 $\uparrow$5.51 85.49 $\uparrow$2.92 24.89 $\uparrow$4.00 45.74 $\uparrow$3.94
Erdos-Renyi 65.99 $\uparrow$4.46 85.37 $\uparrow$2.23 24.54 $\uparrow$3.46 45.45 $\uparrow$3.20
@@
(ImageNet) (ImageNet-R)
Graph prior Acc@1 Acc@5 Acc@1 Acc@5
NACs Scale-Free 75.61 93.93 37.30 54.01
Planted-Partition 75.71 94.09 37.63 54.38
Ring-of-Cliques 76.12 94.35 37.21 53.70
Erdos-Renyi 75.71 93.95 36.48 53.39
[…]
Planted-Partition 76.49 $\uparrow$0.78 94.01 $\downarrow$0.08 38.02 $\uparrow$0.39 53.58 $\downarrow$0.80
Ring-of-Cliques 76.89 $\uparrow$0.77 94.44 $\uparrow$0.09 39.29 $\uparrow$2.08 55.42 $\uparrow$1.72
Erdos-Renyi 76.56 $\uparrow$0.85 94.39 $\uparrow$0.44 38.84 $\uparrow$2.36 54.69 $\uparrow$1.30
Main results. As aforementioned, our teacher model DeiT-Huge and NAC student modules in the first two phases are trained solely on ImageNet. The assembled teacher model achieves a validation accuracy of 81.8%. With 8 A100 80GB GPUs, the training time for all ten student modules in the m2mKD phase is under 12 hours. For E2E training on Tiny-ImageNet, we need to discard a portion of the input tokenizer and the entire output tokenizer in the student modules due to dimension mismatch. Tab. 1 compares the reproduced baselines and our distilled NAC models on Tiny-ImageNet and Tiny-ImageNet-R. The reproduced results slightly outperform the reported values in the original paper. Although the dataset is different from the one used in the m2mKD phase, our distilled models with various graph prior regularizers exhibit an average improvement of 5.3% in IID performance and 3.8% in OOD robustness over the baselines. This indicates that the addition of distilled student modules not only enhances the ability of the final NAC for similar tasks, but also improves its modularity. To ensure reproducibility, we repeat the E2E training phase of the distilled model with scale-free prior three times. The results are presented in Appendix C.
The comparison for ImageNet is shown in Tab. 2. The original paper reports a validation accuracy of 77% for the NAC trained on ImageNet with a scale-free graph prior, while we reproduce the baselines for all four graph priors, achieving a maximum value of 76.1%. Our distilled NACs achieve maximum gains of 1.0% and 2.4% for IID and OOD performance, respectively.
Few-shot. To evaluate the few-shot adaptation performance, we further fine-tune the classifier layer of the distilled NAC with a scale-free prior, which is trained on ImageNet, using a small number of samples from the CIFAR-100 dataset. The hyperparameters and a comprehensive table of results can be found in Appendix B and Appendix D, respectively. We conduct the experiments with 5 different seeds and report the averaged accuracies and corresponding standard deviations in Fig. 6. Our reproduced baselines are at most 10% higher than the original results. It can be observed that the distilled model performs similarly to the baseline with no significant improvement. The standard deviations are relatively large, possibly due to the limited number of repetitions (i.e., the number of seeds). To examine the variation under the same seeds, we rerun the 2-shot experiments once for all five seeds and find that the largest difference reaches around 7% given a fixed seed (see Appendix D for details). Therefore, additional experiments may be necessary to further validate the few-shot performance.
Figure 5: Few-shot performance of NACs on CIFAR-100.
Figure 6: Few-shot performance of DeiT and V-MoE-B models on CIFAR-100 (left) and CUB-100 (right). DI denotes Deep Incubation.
@@ 5 Experiments on MoEs
In a MoE model, some or all of the feedforward networks (FFNs) in a standard monolithic model are replaced with MoE layers. As illustrated in Fig. 4(b), a MoE layer is constructed by multiple experts and a gate, where each expert is essentially an FFN. The gate, typically implemented as a MLP, is responsible for selecting a specified number of experts to process the input. The outputs of the selected experts are then aggregated. The computation performed by a MoE layer can be expressed as [31]:
$y=\sum_{i=1}^{n}G(x)_{i}E_{i}(x)$ (4)
@@ 5.1 Setups
Datasets. We exclusively used the ImageNet-1k dataset for MoE models throughout the pipeline (from preparation to E2E training). The few-shot adaptation ability is evaluated on both CIFAR-100 and CUB-2011 datasets [35].
@@
m2mKD. In this series of experiments, the target model $S$ is set as a Vision MoE Base (V-MoE-B) model [30]. Instead of using the carefully designed gate proposed by [30], we employ the earliest introduced gate in [31] which purely performs a top-$k$ ($k=2$) operation and no additional constraint or balancing loss is applied:
$Gate(x)=Softmax(TopK(W\cdot x)).$ (5)
@@
Unlike the NAC models, the student modules of MoE do not require modification, and all of their learned parameters can be loaded into the target model during the end-to-end training phase. The same as the settings for NAC models, we set the balancing factor $\alpha=0.5$ and softmax temperature $\tau=1.0$. All of the remaining hyperparameters are identical to those used in Deep Incubation for incubating DeiT-B [34] and the student modules are trained for 100 epochs.
@@ 5.2 Results
Main results. We compare m2mKD with three baselines: pure end-to-end training, conventional knowledge distillation, and Deep Incubation. For end-to-end training, we train a V-MoE-B model from scratch for 300 epochs using the same hyperparameters as in the DeiT-B training [34]. In contrast to m2mKD, which performs knowledge distillation at the module level, conventional KD refers to knowledge distillation between complete models. For the sake of fair comparison, we again use DeiT-L as the teacher model and V-MoE-B as the student model. The KL divergence between their output logits is incorporated into the loss with $\alpha=0.5$ and $\tau=1.0$ (see Eq. 3). The training process lasts for 300 epochs. Lastly, we consider the original Deep Incubation approach, where we use their open source checkpoint of the meta model, originally trained for incubating DeiT-B, for the V-MoE-B experiments. The validation results on ImageNet-1k are summarized in Tab. 3. It can be found that the V-MoE-B trained by the pure end-to-end method falls short of its monolithic counterpart, ViT-B, which achieves 81.8% accuracy on ImageNet-1k [34]. This highlights the challenges faced during the training of modular models. Conventional KD underperforms the other methods, with an accuracy 1.9% lower than pure end-to-end training, which is the second lowest. This verifies our assertion that existing knowledge distillation techniques are not necessarily compatible with modular models. On the other hand, the Deep Incubation approach achieves 2.97% higher accuracy than end-to-end training, demonstrating its effectiveness not only for monolithic models, but also for modular models. Our m2mKD approach further increases the accuracy by 0.50%, resulting in the best performance among these methods. Given the remarkable OOD results of NACs in previous experiments, we also investigate the OOD robustness of MoEs on the ImageNet-R dataset, which is rare to be discussed in other MoE-related literatures. Surprisingly, all the MoE models trained using the four different methods achieve near 0% accuracy on the ImageNet-R dataset. Based on these experiments, it appears that NACs are significantly stronger than MoEs in terms of OOD or zero-shot performance.
@@
Training time. In addition to accuracy, Tab. 3 presents the training time ratios relative to the pure end-to-end training approach. m2mKD is slightly slower than Deep Incubation but comparable to pure end-to-end training in terms of time, while conventional KD requires 1.12$\times$ time. Further experiments (Appendix E) demonstrate that the m2mKD ratio can be reduced to 0.57 with marginal performance degradation. For a fair comparison, we exclude the duration of teacher model training in the time calculation for KD and m2mKD methods. Hence, only the time for meta model training, distillation, and end-to-end training phases are considered for m2mKD. It is worth questioning whether the preparation phase of m2mKD can be simplified. Currently, the m2mKD pipeline involves incubating a teacher model. We suppose such a model would be a more knowledgeable teacher for modular students. However, utilizing publicly available pretrained models as the teacher model instead of training a new one from scratch could save considerable time. If this approach is adopted, the total training time would be exactly as stated in the table. We leave the investigation of how teachers trained using different methods influence m2mKD performance in future work.
@@
Few-shot. We evaluate the few-shot adaptation of three models: a DeiT-B model trained using Deep Incubation (DeiT-DI), and two V-MoE-B models trained using Deep Incubation (MoE-DI) and m2mKD (MoE-Ours), respectively. The hyperparameters are listed in Appendix B. During the few-shot training, all parameters in these models are frozen, except for the newly random initialized classifier layer, whose output dimension is set to the number of few-shot classes (8 in all experiments). The experiments for each shot are repeated with five different seeds. The average accuracy for each shot is depicted in Fig. 6. Surprisingly, the DeiT-B model outperforms both V-MoE-B models, which are expected to be capable due to their modularity. This might be attributed to the relatively small dataset size (ImageNet-1k), resulting in insufficient training samples for each expert to learn comprehensively. However, when focusing on the two MoE models, MoE-Ours consistently outperforms MoE-DI by approximately 2% for all shots on CUB-2011, as well as the 1-shot experiment on CIFAR-100. Similar to the few-shot results of NACs, the standard deviations are relatively large, ranging from 2.59 to 12.64. Conducting experiments on a larger dataset is necessary to obtain a more precise evaluation of the few-shot performance.
@@
Ablations. In the proposed m2mKD pipeline, stitch layers are used only during distillation and subsequently discarded to avoid modifying the student model. Although they typically account for a small fraction of parameters in large modular models (e.g., 1% for our V-MoE-B student), their removal may lead to loss of knowledge acquired from the teacher, potentially impacting performance. To investigate this aspect, we conduct experiments preserving stitch layers in the student model. However, the results indicate no noticeable improvement and thus we maintain our decision to discard them. Some additional experiments are carried out based on the setting of preserving stitch layers, including reducing distillation epochs and scaling up. Decreasing the distillation epochs for m2mKD from 100 to 10 results in a 2.1% accuracy drop, yet it remains 2.8% higher than the Deep Incubation baseline. To examine the scalability, we apply m2mKD to V-MoE-Large model and achieve 0.27% higher accuracy on ImageNet-1k compared to Deep Incubation. Given the negligible influence of preserving stitch layers, we believe that the above conclusions remain valid even if they are removed. For further details on these ablation experiments, please refer to Appendix E.